hhanan:

describe the rate of change of this function f(x)=2^x

jhonyy9:

2^1 = 2 2^2 = 4 2^3 = 8 2^4 = 16 2^5 = 32 2^6 = 64 any idea ? @Laylalyssa

hhanan:

times 2

jhonyy9:

yes the next term is always double of the last one

hhanan:

ty

jhonyy9:

np yw